# Posterior Ramifications of Prior Dependence Structures

### v2

@@ 4.2 Practical Implications
Here, we conduct simulations to consider Theorem 2 in practice. To do so, we consider an example adapted from Michimae and Emura (2022) since their recommendations are inconsistent with the results of Theorem 2. They considered a ridge regression model with three regression coefficients in the presence of multicollinearity, where the parameters of the relevant prior copulas were random variables specified using a hierarchical framework. The simplified example for our numerical study adapts aspects of their model for illustrative purposes. Note that we contrast our results with Michimae and Emura’s (2022) findings in Section 4.3.
@@
where $\varepsilon_{i}\sim\mathcal{N}(0,5)$ independently for $i=1,\dots,n$. The assumptions that the linear equation has an intercept of zero and the error terms have known variance reduce the dimensionality of the problem for illustration. That is, $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$. We specify standard normal marginal priors for both $\beta_{1}$ and $\beta_{2}$.
We join these marginal priors with two prior copulas in this numerical study: $c_{1}(\boldsymbol{u})=1$ for $\boldsymbol{u}\in[0,1]^{2}$ corresponds to the independence copula and $c_{2}(\boldsymbol{u})$ corresponds to a two-dimensional $t$-copula with $\nu=4$ degrees of freedom and a diagonal correlation matrix $\boldsymbol{R}$. We select the first copula because it is often assumed that the regression coefficients are independent a priori in Bayesian regression models. The corresponding joint prior for $\beta_{1}$ and $\beta_{2}$ is therefore a standard bivariate normal distribution with diagonal $\boldsymbol{R}$. The choice for the second copula is motivated by $c_{2}(\boldsymbol{u})$ having a local maximum and saddle points to illustrate the results from Theorem 2. Figure 4 visualizes the logarithm of this copula density function. The selected $t$-copula does not accommodate strong negative or positive dependence between $\beta_{1}$ and $\beta_{2}$, but it reflects a greater likelihood of observing extreme values for both $\beta_{1}$ and $\beta_{2}$ relative to their marginal priors. For instance, this might occur if both marginal priors were misspecified.
@@
For each $\boldsymbol{\theta}_{0}$ value, we generated 10000 samples of size $n$ for various sample sizes between 5 and $10^{5}$. Each observation was simulated independently as $y_{i}=\beta_{1,0}x_{i1}+\beta_{2,0}x_{i2}+\varepsilon_{i}$, where $(x_{i1},x_{i2})\sim\mathcal{N}(\boldsymbol{0},I_{2})$ and $\varepsilon_{i}\sim\mathcal{N}(0,5)$ for $i=1,\dots,n$. For each of these 10000 samples, we found both posterior modes $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ and $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ using calculus. We then estimated $Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})$ as the proportion of samples for which $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$, corresponding to the $t$-copula, was closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$. To consider the practical impact of Theorem 2, we also computed the mean absolute difference between $\mathcal{D}_{2}$ and $\mathcal{D}_{1}$ for each sample size $n$. Figure 5 visualizes these results for each of the six scenarios, which we now describe.
@@
Table 2: Empirical coverage of 95% HPD sets for $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$ across 10000 posteriors defined using both prior copulas.
@@
For each case and prior combination, we also estimated empirical coverage as the proportion of the 10000 posteriors for which the 95% HPD set included the parameter value $\boldsymbol{\theta}_{0}$ when $n=\{10^{1},10^{2},10^{3},10^{4},10^{5}\}$. The HPD sets were again estimated using two-dimensional kernel density estimation on posterior draws obtained via sampling-resampling methods. The empirical coverage results are summarized for both prior copulas in Table 2.
@@
Per Table 2, the empirical coverage is similar for both prior copulas at all samples sizes considered in cases 1, 2, 3, and 5. For small sample sizes, the empirical coverage exceeds the nominal value of 95% in cases 1, 2, and 3. Since this trend is observed for both prior copulas, it is caused by informative nature of the relatively well-specified marginal priors for $\beta_{1}$ and $\beta_{2}$. These marginal priors are misspecified in cases 4 and 6, so the empirical coverage is less than 95% for both prior copulas when the sample size is small. The empirical coverage is better when using the independence copula for small $n$ in both cases – even though $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ was generally closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for case 6 when $n$ was small in Figure 5.
4.3 Connections to Previous Contributions
@@
Our simulations demonstrate the value in considering the copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$ when choosing between two prior dependence structures. In particular, we may want to consider the local optima for the copula densities. This numerical study also suggests that copulas have limited ability to reliably prompt more accurate and concentrated posterior distributions around a fixed parameter value $\boldsymbol{\theta}_{0}$. The true value of $\boldsymbol{u}_{0}$ is unknown in practice. And even if $\boldsymbol{u}_{0}$ is such that $c_{2}(\boldsymbol{u}_{0})>c_{1}(\boldsymbol{u}_{0})$, it does not guarantee that $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ will be closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for small or large sample sizes. Regardless of whether the prior dependence structure is a chronically rejected one, choosing a prior copula to improve posterior concentration does not appear to be a sensible objective.
@@
We now contrast our conclusions with Michimae and Emura’s (2022) recommendations. Their numerical studies considered several levels of multicollinearity between the regression covariates for a single $\boldsymbol{\theta}_{0}$ value. For their three regression coefficients, the corresponding $\boldsymbol{u}_{0}$ value lies directly on the upper Fréchet-Hoeffding bound. Their $\boldsymbol{u}_{0}$ value does not necessarily correspond to a local maximum of the Clayton or Gumbel copula density functions they considered. However, the Clayton and Gumbel families of copulas accommodate positive prior dependence and their density functions are generally largest near the upper Fréchet-Hoeffding bound. Their single value is therefore similar to that used in case 2 of our study. They also considered relatively small sample sizes ranging from $n=20$ to 200. Given the results for case 2 in Figure 5, it is not surprising that their normal posterior medians for the regression coefficients were generally closer to their $\boldsymbol{\theta}_{0}$ values when using Clayton and Gumbel copulas. Our numerical results suggest that it is pertinent to consider a variety of $\boldsymbol{\theta}_{0}$ values and sample sizes $n$ before making general statements about the impact of the prior copula on the posterior distribution.
@@
We acknowledge that Michimae and Emura (2022) considered a more complicated model, which leveraged a hierarchical framework to specify the prior copula. Although not the focus of this article, Monte Carlo simulation could likely be used to explore the local optima of such prior copula density functions and extend the results from Theorem 2 to a hierarchical framework. This extension is one of several possible generalizations that could be made to how the analysis prior in (4) is defined.
@@ 5 Discussion
In this paper, we proposed a framework to consider whether prior dependence structures can be retained a posteriori. This framework improves transparency when making inference about posterior dependence structures and helps discard chronically rejected dependence structures for a parameter $\boldsymbol{\theta}$ that cannot be retained as data are observed. Discarding such dependence structures simplifies the prior specification process, particularly when practitioners aim to collect large samples. We also discussed small-sample objectives for prior specification to clarify whether the inability to retain the prior dependence structure presents practical issues for a given posterior analysis. This paper emphasized copula-based priors, but the notion of chronically rejected dependence structures is still applicable to multivariate prior distributions that are not explicitly defined using copulas, such as multivariate normal priors.
@@
Future work could extend the results from this paper to hierarchical settings, in which the hyperparameters of the prior copula are themselves random variables. Moreover, the theoretical results in this paper are based on the standard Bernstein-von Mises theorem. To broaden the applicability of this framework, we could consider nonparametric methods for specifying prior dependence. We could also consider chronically rejected prior dependence structures in the presence of model misspecification (i.e., when $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y})$ that defines the posterior does not coincide with $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ used to generate the data). In that case, we require more complex characterizations of prior dependence than $\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}$ for large samples.

### v3

@@ 4.2 Practical Implications
Here, we conduct simulations to consider Theorem 2 in practice. To do so, we consider an example adapted from Michimae and Emura (Michimae and Emura, 2022) since their recommendations are inconsistent with the results of Theorem 2. They considered a ridge regression model with three regression coefficients in the presence of multicollinearity, where the parameters of the relevant prior copulas were random variables specified using a hierarchical framework. The simplified example for our numerical study adapts aspects of their model for illustrative purposes. Note that we contrast our results with Michimae and Emura’s (Michimae and Emura, 2022) findings in Section 4.3.
@@
where $\varepsilon_{i}\sim\mathcal{N}(0,5)$ independently for $i=1,\dots,n$. The assumptions that the linear equation has an intercept of zero and the error terms have known variance reduce the dimensionality of the problem for illustration. That is, $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$. We specify standard normal marginal priors for both $\beta_{1}$ and $\beta_{2}$. We join these marginal priors with two prior copulas in this numerical study: $c_{1}(\boldsymbol{u})=1$ for $\boldsymbol{u}\in[0,1]^{2}$ corresponds to the independence copula and $c_{2}(\boldsymbol{u})$ corresponds to a two-dimensional $t$-copula with $\nu=4$ degrees of freedom and a diagonal correlation matrix $\boldsymbol{R}$ (i.e, the $2\times 2$ identity matrix). Using the notation from Corollary 1, both copulas are such that $\tau^{p}_{\beta_{1},\beta_{2}}=0$.
We select the first copula because it is often assumed that the regression coefficients are independent a priori in Bayesian regression models. The corresponding joint prior for $\beta_{1}$ and $\beta_{2}$ is therefore a standard bivariate normal distribution with diagonal $\boldsymbol{R}$. The choice for the second copula is motivated by $c_{2}(\boldsymbol{u})$ having a local maximum and saddle points to illustrate the results from Theorem 2. Figure 4 visualizes the logarithm of this copula density function. The selected $t$-copula does not accommodate strong negative or positive dependence between $\beta_{1}$ and $\beta_{2}$, but it reflects a greater likelihood of observing extreme values for both $\beta_{1}$ and $\beta_{2}$ relative to their marginal priors. For instance, this might occur if both marginal priors were misspecified.
@@
For each $\boldsymbol{\theta}_{0}$ value, we generated 10000 samples of size $n$ for various sample sizes between 5 and $10^{5}$. Each observation was simulated independently as $y_{i}=\beta_{1,0}x_{i1}+\beta_{2,0}x_{i2}+\varepsilon_{i}$, where $(x_{i1},x_{i2})\sim\mathcal{N}(\boldsymbol{0},I_{2})$ and $\varepsilon_{i}\sim\mathcal{N}(0,5)$ for $i=1,\dots,n$. This data generation process implies $\tau(\boldsymbol{\theta}_{0})_{\beta_{1},\beta_{2}}=0$. For each of these 10000 samples, we found both posterior modes $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ and $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ by numerically finding the zeros of the partial derivatives. We then estimated $Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})$ as the proportion of samples for which $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$, corresponding to the $t$-copula, was closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$. To consider the practical impact of Theorem 2, we also computed the mean absolute difference between $\mathcal{D}_{2}$ and $\mathcal{D}_{1}$ for each sample size $n$. Figure 5 visualizes these results for each of the six scenarios, which we now describe.
@@
Table 2: Empirical coverage of 95% HPD sets for $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$ across 10000 posteriors defined using both prior copulas. The median area of the HPD set across the 10000 posteriors for each case and prior combination is given in parentheses.
@@
For each case and prior combination, we also estimated empirical coverage as the proportion of the 10000 posteriors for which the 95% HPD set included the parameter value $\boldsymbol{\theta}_{0}$ when $n=\{10^{1},10^{2},10^{3},10^{4},10^{5}\}$. The HPD sets were again estimated using two-dimensional kernel density estimation on posterior draws obtained via sampling-resampling methods. We estimated the median area of the 10000 HPD sets for each case and prior combination using quasi-Monte Carlo methods (Lemieux, 2009). The empirical coverage and median area results are summarized for both prior copulas in Table 2.
@@
Per Table 2, the empirical coverage is similar for both prior copulas at all samples sizes considered in cases 1, 2, 3, and 5. For small sample sizes, the empirical coverage exceeds the nominal value of 95% in cases 1, 2, and 3. Since this trend is observed for both prior copulas, it is caused by the informative nature of the relatively well-specified marginal priors for $\beta_{1}$ and $\beta_{2}$. These marginal priors are misspecified in cases 4 and 6, so the empirical coverage is less than 95% for both prior copulas when the sample size is small. The empirical coverage is better when using the independence copula for small $n$ in both cases – even though $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ was generally closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for case 6 when $n$ was small in Figure 5.
Neither prior copula reliably prompts greater posterior concentration and coverage. The $t$-copula gives rise to slightly smaller median areas of the 95% HPD sets for cases 1, 2, 3, and 5 when $n=10^{1}$ or $10^{2}$. However, the independence copula yields smaller median areas at these sample sizes for case 4. In case 6, the median area is much smaller for the $t$-copula when $n=10^{1}$, but this increase in posterior concentration is accompanied by a decrease in coverage. The independence copula prompts both greater coverage and posterior concentration for case 6 when $n=10^{2}$.
4.3 Connections to Other Work
@@
Our simulations demonstrate the value in considering the copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$ when choosing between two prior dependence structures. In particular, we may want to consider the local optima for the copula densities. This numerical study also suggests that copulas have limited ability to reliably prompt more accurate and concentrated posterior distributions around a fixed parameter value $\boldsymbol{\theta}_{0}$. The $t$-copula only reliably prompted a more accurate and concentrated posterior when $\boldsymbol{u}_{0}$ was in a neighbourhood of (0.5, 0.5). To reap these benefits, we would need to know that $\boldsymbol{u}_{0}$ is in a particular region of $[0,1]^{d}$ a priori. In that event, we should likely make our marginal priors more informative to reflect this knowledge. Such strategies would be more related to past work on model identifiability (Gelfand and Sahu, 1999; Eberly and Carlin, 2000; Gustafson, 2005).
In practice, the true value of $\boldsymbol{u}_{0}$ is instead unknown. And even if $\boldsymbol{u}_{0}$ is such that $c_{2}(\boldsymbol{u}_{0})>c_{1}(\boldsymbol{u}_{0})$, our numerical studies demonstrate that $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ is not guaranteed to be closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for small or large sample sizes. Both copula priors in our numerical study were such that $\tau^{p}_{\beta_{1},\beta_{2}}=\tau(\boldsymbol{\theta}_{0})=0$; neither prior dependence structure satisfied the conditions for Corollary 1. Regardless of whether the prior dependence structure is a chronically rejected one, choosing a prior copula to improve posterior concentration is likely not a sensible objective.
@@
We now contrast our conclusions with Michimae and Emura’s (Michimae and Emura, 2022) recommendations. Their numerical studies considered several levels of multicollinearity between the regression covariates for a single $\boldsymbol{\theta}_{0}$ value. The joint prior for their three regression coefficients leveraged a vine structure with three bivariate copulas. For each bivariate copula, the corresponding $\boldsymbol{u}_{0}$ value lies directly on the upper Fréchet-Hoeffding bound. Each $\boldsymbol{u}_{0}$ value does not necessarily correspond to a local maximum of the Clayton or Gumbel copula density functions they considered. However, the Clayton and Gumbel families of copulas accommodate positive prior dependence and their density functions are generally largest near the upper Fréchet-Hoeffding bound. Their single $\boldsymbol{\theta}_{0}$ value is therefore similar to that used in case 2 of our study. They also considered relatively small sample sizes ranging from $n=20$ to 200. Given the results for case 2 in Figure 5, it is not surprising that their normal posterior medians for the regression coefficients were generally closer to their $\boldsymbol{\theta}_{0}$ values when using Clayton and Gumbel copulas, despite this result not holding more generally. Our numerical results suggest that it is pertinent to consider a variety of $\boldsymbol{\theta}_{0}$ values and sample sizes $n$ before making general statements about the impact of the prior copula on the posterior distribution.
@@
We acknowledge that Michimae and Emura (Michimae and Emura, 2022) considered a more complicated model, which leveraged a hierarchical framework to specify the prior copula. Although not the focus of this article, Monte Carlo simulation could likely be used to explore the local optima of such prior copula density functions and extend the results from Theorem 2 to a hierarchical framework. This extension is one of several possible generalizations that could be made to the definition of the analysis prior in (4).
@@ 5 Discussion
In this paper, we proposed a framework to consider whether prior dependence structures can be retained a posteriori. This framework informs conversations between statisticians and practitioners to promote alignment between the flexibility in the prior dependence structure and the objectives for posterior analysis. Discarding chronically rejected dependence structures that cannot be retained as data are observed simplifies the prior specification process when practitioners aim to collect large samples or prioritize inference on the dependence structure. We also discussed small-sample objectives for prior specification to clarify that the inability to retain the prior dependence structure may not present practical issues for various posterior analyses. This paper emphasized copula-based priors, but the notion of chronically rejected dependence structures is still applicable to multivariate prior distributions that are not explicitly defined using copulas, such as multivariate normal priors.
@@
Future work could extend the results from this paper to hierarchical settings, in which the hyperparameters of the prior copula are themselves random variables. Moreover, the theoretical results in this paper are based on the standard Bernstein-von Mises theorem. To broaden the applicability of this framework, we could consider nonparametric methods for specifying prior dependence. We could also consider chronically rejected prior dependence structures in the presence of model misspecification (i.e., when $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y})$ that defines the posterior does not coincide with $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ used to generate the data). In that case, we require more complex characterizations of prior dependence than $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ for large samples.